# Physics Informed Kolmogorov-Arnold Neural Networks for Dynamical Analysis via Efficent-KAN and WAV-KAN

### v1

@@ V.1 System of Linear Ordinary Differential Equations
We have utilized the efficient-KAN method to solve the coupled-linear differential equation (V.1), employing collocation points $N_{x}=100$ equally distributed on the interval $0\leq x\leq 1$ [35]. Note that we have observed that there is not much difference between the performance of WAV-KAN and efficient-KAN for this problem, but the convergence of WAV-KAN is faster than efficient-KAN. The model was trained for 10000 epochs and AdamW optimizer is used with a learning rate $\eta=0.001$. At the end of the 3000 epochs, the recorded loss was of the order of $10^{-4}.$ Further details about the model used are listed in Table 3. The comparison of the PIKAN predicted solution and the numerically exact solution of $u(x)$ and $v(x)$ is shown in Fig 3(b).
@@ V.2 System of Nonlinear Ordinary Differential Equations
We have utilized the WAV-KAN method to solve the coupled-nonlinear differential equation (18), employing collocation points $N_{x}=100$ equally distributed on the interval $0\leq x\leq 1$ [35]. The interesting point is that there is not much difference between the performance of WAV-KAN and efficient-KAN for this particular problem but the convergence rate of WAV-KAN was faster than efficient-KAN. The model was trained for 3000 epochs and AdamW optimizer is used with a learning rate $\eta=0.001$. At the end of the 3000 epochs, the recorded loss was of the order of $10^{-4}.$ Further details about the model used is listed in Table 4. The comparison of PIKAN predicted solution and the numerically computed solution is shown in Fig 4(b).
Figure 4: (a) Illustrates the relationship between loss (21) and epoch, (b) A comparison between PIKAN predicted solution and numerically exact sollution of Eq (18).
Note that in Fig 4(b), the PIKAN solution for $\hat{u}(x)$ does not precisely match the exact numerical solution for $u(x)$. This deviation is observed when a data-free approach is employed. This is because the data-free approaches often struggle with convergence, especially for highly nonlinear differential equations leading to chaotic dynamics. Even if they converge, they might not accurately capture the system’s details, leading to significant errors. To achieve an exact fit in such cases, a data-driven approach is beneficial as it guides the model towards the correct solution.
@@ VIII Comparing Physics-Informed Models: PINN and PIKAN
Let us now explore how PIKAN simplifies hyperparameter tuning compared to PINNs. First, for the case of the system of two linear equations (V.1), two separate neural networks have been trained in [35]: The PINN model approximating $u(x)$ [Eq. (V.1)] has 20 neurons and 4 hidden layers, and the PINN model for $v(x)$ has 10 neurons and 5 hidden layers. A single PIKAN model presented in Section V.1 with architecture $[1,2,3,2]$ (recall that there is only two intermediate layers) already performed well.

### v2

@@ V.1 System of Linear Ordinary Differential Equations
We have utilized the efficient-KAN method to solve the coupled-linear differential equation (V.1), employing collocation points $N_{x}=100$ equally distributed on the interval $0\leq x\leq 1$ [35]. Note that we have observed that there is not much difference between the performance of WAV-KAN and efficient-KAN for this problem, but the convergence of WAV-KAN is faster than efficient-KAN. The model was trained for 10000 epochs and AdamW optimizer is used with a learning rate $\eta=0.001$. At the end of the 10000 epochs, the recorded loss was of the order of $10^{-5}.$ Further details about the model used are listed in Table 3. The comparison of the PIKAN predicted solution and the numerically exact solution of $u(x)$ and $v(x)$ is shown in Fig 3(b).
@@ V.2 System of Nonlinear Ordinary Differential Equations
We have utilized the WAV-KAN method to solve the coupled-nonlinear differential equation (18), employing collocation points $N_{x}=100$ equally distributed on the interval $0\leq x\leq 1$ [35]. The interesting point is that there is not much difference between the performance of WAV-KAN and efficient-KAN for this particular problem but the convergence rate of WAV-KAN was faster than efficient-KAN. The model was trained for 10000 epochs and AdamW optimizer is used with a variable learning rate starting from $\eta=0.01$, and at the end of every 1000 epochs the learning rate is updated following the given rule, $\eta_{new}=\frac{\eta_{old}}{5}$ At the end of the 10000 epochs, the recorded loss was of the order of $10^{-5}.$ Further details about the model used is listed in Table 4. The comparison of PIKAN predicted solution and the numerically computed solution is shown in Fig 4(b).
Figure 4: (a) Illustrates the relationship between loss (21) and epoch, (b) A comparison between PIKAN predicted solution and numerically exact solution of Eq (18).
@@ VIII Comparing Physics-Informed Models: PINN and PIKAN
Let us now explore how PIKAN simplifies hyperparameter tuning compared to PINNs. First, for the case of the system of two linear equations (V.1), two separate neural networks have been trained in [35]: The PINN model approximating $u(x)$ [Eq. (V.1)] has 20 neurons and 4 hidden layers, and the PINN model for $v(x)$ has 10 neurons and 5 hidden layers. A single PIKAN model presented in Section V.1 with architecture $[1,2,3,2]$ (recall that there is only two intermediate layers) already performed well. For the system of equations represented by (18), the PINN model required 8 hidden layers with 8 neurons each, as detailed in [35]. In contrast, the corresponding PIKAN model (in Section V.2) achieved similar accuracy with just a single layer having an architecture of $[1,7,2]$.